# Shortcuts to adiabaticity for coherent atom transport in an adjustable family of two-dimensional optical lattices

### v1

@@ III.1 Invariant-based inverse engineering and the Lewis-Leach family of Hamiltonians
$V(x,t)=-F(t)x+\frac{m}{2}\>\omega^{2}(t)x^{2}+\frac{1}{\rho^{2}(t)}U\left(\frac{x-\alpha(t)}{\rho(t)}\right)\>,$ (15)
where $\rho(t)$, $\alpha(t)$, $\omega(t)$, and $F(t)$ are arbitrary functions of time satisfying certain auxiliary equations Tor . In particular, the functions $\rho(t)$ and $\omega(t)$ ought to satisfy the nonlinear second-order ordinary differential equation
@@
which is known as the Ermakov equation Erm , where $\omega_{0}$ is a constant. At the same time, the functions $\alpha(t)$, $\omega(t)$, and $F(t)$ have to obey the condition
$\ddot{\alpha}+\omega^{2}(t)\alpha=\frac{F(t)}{m}\>.$ (17)
@@
$I(t)$ $=$ $\frac{1}{2m}\big[\rho(p-m\dot{\alpha})-m\dot{\rho}(x-\alpha)]^{2}$
$+$ $\frac{m}{2}\>\omega^{2}_{0}\left(\frac{x-\alpha}{\rho}\right)^{2}+U\left(\frac{x-\alpha}{\rho}\right)\>.$
@@ IV Dynamical-lattice trajectories
In what follows, we apply the general STA results reviewed in Sec. III to coherent atom transport in the adjustable family of 2D optical lattices under consideration. We first utilize the STA approach to obtain the classical path of the potential minima in a moving lattice. We do so within the framework of a single-atom Hamiltonian with the harmonically approximated potential of Eq. (7).
In the atom-transport problem under consideration, the $x$- and $y$ components of the (time-dependent) dynamical-lattice trajectory vector will be denoted by $q_{0}^{\mathrm{x}}(\bm{\lambda}_{\mathrm{x}};t)$ and $q_{0}^{\mathrm{y}}(\bm{\lambda}_{\mathrm{y}};t)$. Here $\bm{\lambda}=\left(\bm{\lambda}_{\mathrm{x}},\bm{\lambda}_{\mathrm{y}}\right)^{\intercal}$ is the (real-valued) control vector, where $\bm{\lambda}_{\mathrm{x}}\equiv[\lambda^{(1)}_{\mathrm{x}},\ldots,\lambda^{(6)}_{\mathrm{x}}$] and $\bm{\lambda}_{\mathrm{y}}\equiv[\lambda^{(1)}_{\mathrm{y}},\ldots,\lambda^{(6)}_{\mathrm{y}}$] are the reduced control vectors corresponding to the displacements in the $x$- and $y$ directions, respectively. The latter are assumed to satisfy the following conditions:
$q_{0}^{\mathrm{x}}(\bm{\lambda}_{\mathrm{x}};\>jt_{f}/7)$ $=$ $\lambda^{(j)}_{\mathrm{x}}\quad(\>j=1,\ldots,6\>)\>,$
$q_{0}^{\mathrm{y}}(\bm{\lambda}_{\mathrm{y}};\>jt_{f}/7)$ $=$ $\lambda^{(j)}_{\mathrm{y}}\>$. (23)
The STA dynamical-lattice trajectory vector, whose $x$- and $y$ components are a special case of $q_{0}^{\mathrm{x}}(\bm{\lambda}_{\mathrm{x}};t)$ and $q_{0}^{\mathrm{y}}(\bm{\lambda}_{\mathrm{y}};t)$, corresponds to the control vector $\bm{\lambda}_{0}\equiv\left(\bm{\lambda}_{\mathrm{x},0},\bm{\lambda}_{\mathrm{y},0}\right)^{\intercal}$, with components $\lambda^{(j)}_{\mathrm{x},0}$ and $\lambda^{(j)}_{\mathrm{y},0}$ $(j=1,\ldots,6)$. Thus, the STA trajectory is described by $q_{0}^{\mathrm{x}}(\bm{\lambda}_{\mathrm{x},0};\>t)$ and $q_{0}^{\mathrm{y}}(\bm{\lambda}_{\mathrm{y},0};\>t)$.
@@
$F_{\mathrm{x}}(t)=m\left[\omega^{2}_{\mathrm{x}}q_{0,x}(t)-a_{\mathrm{x}}\right]\>.$ (26)
@@
$F_{\mathrm{y}}(t)=m\omega^{2}_{\mathrm{y}}q_{0,y}(t)\>.$ (27)
@@
One peculiar aspect of these last auxiliary equations is the presence of $a_{x}$ on the right-hand-side of Eq. (28), resulting from the terms linear in $x$ in the approximated potentials $V_{\mathrm{L}}$ [cf. Eq. (7)]. The presence of $a_{x}$ renders the problem at hand analogous to transport problems involving a constant force (e.g., gravity) Tor . It should be noted that Eq. (28) can be reduced to the form of Eq. (29) by a slight redefinition of $q_{\mathrm{c}}^{\mathrm{x}}(t)$ – namely, by introducing $q_{\mathrm{c}}^{\mathrm{x}}(t)\equiv q_{\mathrm{c}}^{\mathrm{x}}(t)+a_{x}/\omega_{\mathrm{x}}^{2}$. In this manner, the problem of atom transport in the $x$ direction, equally like the one in the $y$-direction, is reduced to a well-known type of atom-transport problems solvable via inverse engineering Tor .
Figure 4: Path of a potential minimum as a function of time, obtained using the STA approach, for transport times $t_{f}$ comparable to (a), and an order of magnitude longer than (b) the internal timescale $T_{\mathrm{x}}$.
@@
$q_{0,u}(t)=d_{u}\sum_{n=3}^{9}b_{n,u}\left(\frac{t}{t_{\mathrm{f}}}\right)^{n}\quad(u=x,y)\>,$ (30)
@@
$b_{8,u}$ $=$ $-315\quad,\quad b_{9,u}=70\>.$
The obtained solution for the STA dynamical-lattice trajectory – more precisely, the $x$ component $q_{0,x}$ of the path of the potential minimum – is shown as a function of the transport time $t_{\mathrm{f}}$ in Fig. 4.
@@ V Calculation of atomic dynamics
In what follows, we briefly present our chosen approach for evaluating the single-atom dynamics corresponding to STA-based dynamical-lattice trajectories [cf. Sec. IV]. This approach is based on the use of the FSOM for numerically solving the relevant TDSE (for a short introduction into this method, see Appendix A).
The crucial prerequisite for computing single-atom dynamics in the problem at hand amounts to evaluating the ground-state wave-function of the relevant optical-lattice potential. Within the general framework of the atom-transport problem (cf. Sec. II.1), that wave function represents the initial ($t=0$) atomic wave packet, i.e. the initial condition for single-atom transport. This wave function can be obtained to requisite accuracy using the plane-wave-expansion method (for details, see, e.g., Sto (b)), which entails Fourier expansions of the underlying lattice potential and the periodic part of single-particle Bloch wave functions in terms of reciprocal-lattice vectors. Yet, the same can be accomplished by employing the FSOM in conjunction with the ITE approach (for details, see Appendix B). Given that in this study we employ the FSOM for the dynamics part, this last scheme appears as the more convenient choice.
@@
where $\delta\dot{\mathbf{q}}_{0}(t)\equiv\dot{\mathbf{q}}_{0}(t+\delta t)-\dot{\mathbf{q}}_{0}(t)$. It is worthwhile to mention that two additional multiplicative terms – that do not depend on the momentum- and coordinate operators, hence leading to irrelevant time-dependent global phase factors – have been dropped on the RHS of the last equation, analogously to the dropped terms in Eq. (33) above.
With the aid of an analog of Eq. (A) and by performing a Fourier transformation of the $\hat{\mathbf{p}}-$dependent terms in (37), switching also to the coordinate representation of the momentum operator [i.e. $\hat{\mathbf{p}}\rightarrow(\hbar/i)\nabla$], we arrive at the FSOM-based (second-order) time-stepping scheme for the atom-transport problem at hand:
@@
In order to alleviate the computational burden in the problem at hand, we exploit the discrete translational symmetry of the system and restrict ourselves to the displacement of one unit cell of the original optical lattice. The appropriate choice of this unit cell has to be consistent with the transport direction. For instance, in order to evaluate transport in the $x$- and $y$ directions we make use of a rectangular unit cell (indicated by white dashed lines in Fig. 2).
@@
The spatial Fourier transformation is carried out using the FFT algorithm (cf. Appendix ANRc . To be more specific, we make use of a discrete three-dimensional $N_{x}\times N_{y}\times N_{z}$ grid with $N_{x}=200$, $N_{y}=300$, and $N_{z}=100$ points in the computational window whose dimensions (in units of the respective characteristic lenghtscales $l_{\mathrm{x}}$, $l_{\mathrm{y}}$, and $l_{\mathrm{z}}$) are $100$, 100100, and $500$, respectively. For the propagation of the TDSE in the comoving frame [cf. Eq. (33)], we make use of the adaptive approach NRc with a maximal relative error of $10^{-4}$; the latter could – strictly speaking – require additional (smaller in size) substeps in time. Yet, it turned out to be sufficient to take $N_{t}$ between $20$ and $100$ in all our runs.

### v2

@@ III.1 Invariant-based inverse engineering and the Lewis-Leach family of Hamiltonians
$V(x,t)=-F(t)x+\frac{m}{2}\>\omega^{2}(t)x^{2}+\frac{1}{\rho^{2}(t)}U\left(\frac{x-\beta(t)}{\rho(t)}\right)\>,$ (17)
where $\rho(t)$, $\beta(t)$, $\omega(t)$, and $F(t)$ are arbitrary functions of time satisfying certain auxiliary equations Tor . In particular, the functions $\rho(t)$ and $\omega(t)$ ought to satisfy the nonlinear second-order ordinary differential equation
@@
which is known as the Ermakov equation Erm , where $\omega_{0}$ is a constant. At the same time, the functions $\beta(t)$, $\omega(t)$, and $F(t)$ have to obey the condition
$\ddot{\beta}+\omega^{2}(t)\beta=\frac{F(t)}{m}\>.$ (19)
@@
$I(t)$ $=$ $\frac{1}{2m}\big[\rho(p-m\dot{\beta})-m\dot{\rho}(x-\beta)]^{2}$
$+$ $\frac{m}{2}\>\omega^{2}_{0}\left(\frac{x-\beta}{\rho}\right)^{2}+U\left(\frac{x-\beta}{\rho}\right)\>.$
@@ III.2 Harmonic transport: STA treatment
III.3 Basics of the eSTA method
While STA-based control protocols have established themselves in a variety of quantum systems STA , their modification – which became known as eSTA – has been proposed Whi (a). The motivation behind this enhanced version of the STA approach, which is inspired in part by optimal-control methods Whi (a), is to facilitate the development of efficient control schemes for systems not directly amenable to an STA-type treatment. The principal idea of eSTA is to approximate the original Hamiltonian of such a system by a simpler one that admits a straightforward design of an STA-based protocol. Assuming that this last protocol is nearly optimal even when applied to the original Hamiltonian, its eSTA counterpart is constructed using a gradient expansion in the control-parameter space. This crucial assumption underscores the semi-heuristic character of the eSTA method, which – in principle – is not guaranteed to outperform STA Whi (a). Yet, it has already been demonstrated that eSTA is superior to STA in several classes of quantum-control problems HauSto ; Odell (1).
The initial step that the application of the eSTA method to a system described by the Hamiltonian $H_{\textrm{S}}$ entails amounts to obtaining an STA solution for a “close,” approximated Hamiltonian $H_{0}$; the latter corresponds to the value $\bm{\zeta}_{0}$ of the vector parameter $\bm{\zeta}\in\mathbb{R}^{n}$ Whi (a). It is assumed that a value $\mu_{\textrm{S}}$ of a scalar parameter $\mu$ can be found such that a series expansion
$H_{\textrm{S}}=\sum_{k=0}^{\infty}\mu_{\textrm{S}}^{k}\,H^{(k)}\>,$ (25)
has $H_{0}$ as its zeroth-order term [i.e. $H^{(0)}\equiv H_{0}$].
In order to find an eSTA solution for $H_{\textrm{S}}$ based on a previously found STA solution for $H_{0}$, the control vector $\bm{\zeta}_{\mathrm{S}}$ corresponding to $H_{\textrm{S}}$ can be written as a sum of the STA control vector $\bm{\zeta}_{0}$ and an auxiliary control vector $\bm{\alpha}$, i.e. $\bm{\zeta}_{\mathrm{S}}=\bm{\zeta}_{0}+\bm{\alpha}$. The value of $\bm{\alpha}$ that corresponds to the desired eSTA solution, will be denoted by $\bm{\epsilon}$ in the following.
As stated above, the principal assumption underlying the eSTA scheme is that the STA-based protocol corresponding to the Hamiltonian $H_{0}$ is near-optimal even when applied to the full Hamiltonian $H_{\textrm{S}}$ Whi (a). Another important assumption within the eSTA approach is that the deviation of the fidelity $\mathcal{F}$ from its maximal value is quadratic in the difference $\alpha-\epsilon$, i.e. the fidelity can be written in the form Whi (a)
$\mathcal{F}\left(\mu_{\textrm{S}},\bm{\zeta}_{0}+\alpha\frac{\bm{\nabla}\mathcal{F}(\mu_{\textrm{S}},\bm{\zeta}_{0})}{\rVert\bm{\nabla}\mathcal{F}(\mu_{\textrm{S}},\bm{\zeta}_{0})\rVert}\right)\approx 1-c\left(\alpha-\epsilon\right)^{2}\>,$ (26)
where $\epsilon\equiv\rVert\bm{\epsilon}\rVert$, $\alpha\equiv\rVert\bm{\alpha}\rVert$, and $c$ is a positive constant.
The above assumptions and a Taylor expansion of the left-hand-side of Eq. (26) around $\epsilon=\alpha$, straightforwardly lead to  Whi (a)
$\bm{\epsilon}\approx\frac{2\left[1-\mathcal{F}(\mu_{\textrm{S}},\bm{\zeta}_{\textrm{S}})\right]\bm{\nabla}\mathcal{F}(\mu_{\textrm{S}},\bm{\zeta}_{0})}{\rVert\bm{\nabla}\mathcal{F}(\mu_{\textrm{S}},\bm{\zeta}_{0})\rVert^{2}}\>.$ (27)
To second order in $\mu_{\textrm{S}}$ the fidelity is given by Whi (a)
$\mathcal{F}(\mu_{\textrm{S}},\bm{\zeta}_{\textrm{S}})\approx 1-\frac{1}{\hbar^{2}}\sum_{n=1}^{\infty}|G_{n}|^{2}\>,$ (28)
where $G_{n}$ is an auxiliary scalar function, given in terms of the transport modes of $H_{0}$ [cf. Sec. III.1] by
$G_{n}=\int_{0}^{t_{f}}dt\braket{\Psi_{n}(t)|\left[H_{\textrm{S}}(\bm{\zeta}_{0};t)-H_{0}(\bm{\zeta}_{0};t)\right]|\Psi_{0}(t)}\>.$ (29)
Likewise, up to second order in $\mu_{\textrm{S}}$ an analogous expression for the gradient of $\mathcal{F}(\mu_{\textrm{S}},\bm{\zeta}_{0})$ reads Whi (a)
$\bm{\nabla}\mathcal{F}(\mu_{\textrm{S}},\bm{\zeta}_{0})\approx-\frac{2}{\hbar^{2}}\sum_{n=1}^{\infty}\textrm{Re}\left(G_{n}\>\bm{K}_{n}^{*}\right)\>,$ (30)
with $\bm{K}_{n}$ being an auxiliary vector function:
$\bm{K}_{n}=\int_{0}^{t_{f}}dt\braket{\Psi_{n}(t)|\nabla_{\zeta}H_{\textrm{S}}(\bm{\zeta};t)\big.|{}_{\bm{\zeta}=\bm{\zeta}_{0}}|\Psi_{0}(t)}\>.$ (31)
It should be stressed at this point that the evaluation of these two auxiliary functions ($G_{n}$ and $\bm{K}_{n}$) requires only the knowledge of the full Hamiltonian $H_{\textrm{S}}(\bm{\zeta};t)$ of the system and the known STA solutions for the idealized system governed by Hamiltonian $H_{0}$.
In terms of $G_{n}$ and $\bm{K}_{n}$, the eSTA correction vector $\bm{\epsilon}$ can be expressed as
$\bm{\epsilon}=-\frac{\left(\sum_{n=1}^{N}|G_{n}|^{2}\right)\sum_{n=1}^{N}\textrm{Re}\left(G_{n}^{\ast}\bm{K}_{n}\right)}{\left\rVert\sum_{n=1}^{N}\textrm{Re}\left(G_{n}^{\ast}\bm{K}_{n}\right)\right\rVert^{2}}\>,$ (32)
where the infinite sum over $n$ is truncated to the first $N$ terms (i.e. $N$ serves as the cut-off parameter). From this last expression, $\bm{\epsilon}$ can straightforwardly be computed numerically provided that $G_{n}$ and $\bm{K}_{n}$ are previously calculated using Eqs. (29) and (31), respectively.
A general heuristic argument can be made that provides an explanation as to why eSTA-based control protocols should be more robust (i.e. less sensitive) to parameter deviations than their STA-based counterparts. Namely, from the very assumption – inherent in the construction of the eSTA approach – that the state fidelity behaves as a quadratic function in the vicinity of its maximum [cf. Eq. (26)] and the fact that the fidelity corresponding to an eSTA scheme is higher than that of the attendant STA scheme, it follows that the derivative of the eSTA fidelity curve should be smaller than that of its STA counterpart. This, in turn, signifies increased robustness of the eSTA scheme.
@@ IV Dynamical-lattice trajectories
In what follows, we apply the general STA- and eSTA results reviewed in Sec. III to coherent atom transport in the adjustable family of 2D optical lattices under consideration. We first utilize the STA approach to obtain the classical path of the potential minima in a moving lattice (see Sec. IV.1 below). We do so within the framework of a single-atom Hamiltonian with the harmonically approximated potential of Eq. (9). We then address the same problem using the more sophisticated eSTA approach (see Sec. IV.2 below), which takes into account the full, anharmonic optical-lattice potential.
In the atom-transport problem under consideration, the $x$- and $y$ components of the (time-dependent) dynamical-lattice trajectory vector will be denoted by $q_{0}^{\mathrm{x}}(\bm{\zeta}_{\mathrm{x}};t)$ and $q_{0}^{\mathrm{y}}(\bm{\zeta}_{\mathrm{y}};t)$. Here $\bm{\zeta}=\left(\bm{\zeta}_{\mathrm{x}},\bm{\zeta}_{\mathrm{y}}\right)^{\intercal}$ is the (real-valued) control vector, where $\bm{\zeta}_{\mathrm{x}}\equiv[\zeta^{(1)}_{\mathrm{x}},\ldots,\zeta^{(6)}_{\mathrm{x}}$] and $\bm{\zeta}_{\mathrm{y}}\equiv[\zeta^{(1)}_{\mathrm{y}},\ldots,\zeta^{(6)}_{\mathrm{y}}$] are the reduced control vectors corresponding to the displacements in the $x$- and $y$ directions, respectively. The latter are assumed to satisfy the following conditions:
$q_{0}^{\mathrm{x}}(\bm{\zeta}_{\mathrm{x}};\>jt_{f}/7)$ $=$ $\zeta^{(j)}_{\mathrm{x}}\quad(\>j=1,\ldots,6\>)\>,$
$q_{0}^{\mathrm{y}}(\bm{\zeta}_{\mathrm{y}};\>jt_{f}/7)$ $=$ $\zeta^{(j)}_{\mathrm{y}}\>$. (33)
IV.1 STA dynamical-lattice trajectories
The STA dynamical-lattice trajectory vector, whose $x$- and $y$ components are a special case of $q_{0}^{\mathrm{x}}(\bm{\zeta}_{\mathrm{x}};t)$ and $q_{0}^{\mathrm{y}}(\bm{\zeta}_{\mathrm{y}};t)$, corresponds to the control vector $\bm{\zeta}_{0}\equiv\left(\bm{\zeta}_{\mathrm{x},0},\bm{\zeta}_{\mathrm{y},0}\right)^{\intercal}$, with components $\zeta^{(j)}_{\mathrm{x},0}$ and $\zeta^{(j)}_{\mathrm{y},0}$ $(j=1,\ldots,6)$. Thus, the STA trajectory is described by $q_{0}^{\mathrm{x}}(\bm{\zeta}_{\mathrm{x},0};\>t)$ and $q_{0}^{\mathrm{y}}(\bm{\zeta}_{\mathrm{y},0};\>t)$.
@@ IV.1 STA dynamical-lattice trajectories
$F_{\mathrm{x}}(t)=m\left[\omega^{2}_{\mathrm{x}}q_{0}^{\mathrm{x}}(t)-a_{\mathrm{x}}\right]\>.$ (36)
@@
$F_{\mathrm{y}}(t)=m\omega^{2}_{\mathrm{y}}q_{0}^{\mathrm{y}}(t)\>.$ (37)
@@
One peculiar aspect of these last auxiliary equations is the presence of $a_{\mathrm{x}}$ on the RHS of Eq. (38), resulting from the terms linear in $x$ in the approximated potentials $V_{\mathrm{L}}$ [cf. Eq. (9)]. The presence of $a_{\mathrm{x}}$ renders the problem at hand analogous to transport problems involving a constant force (e.g., gravity) Tor . It should be noted that Eq. (38) can be reduced to the form of Eq. (39) by a slight redefinition of $q_{\mathrm{c}}^{\mathrm{x}}(t)$ – namely, by introducing $q_{\mathrm{c}}^{\mathrm{x}}(t)\equiv q_{\mathrm{c}}^{\mathrm{x}}(t)+a_{\mathrm{x}}/\omega_{\mathrm{x}}^{2}$. In this manner, the problem of atom transport in the $x$ direction, equally like the one in the $y$-direction, is reduced to a well-known type of atom-transport problems solvable via inverse engineering Tor .
Figure 4: (Color online) Path of a potential minimum as a function of time, obtained using the STA approach, for transport times $t_{f}$ comparable to (a), and an order of magnitude longer than (b) the internal timescale $T_{\mathrm{x}}$.
@@
$q_{0}^{u}(t)=d_{u}\sum_{n=3}^{9}b_{n,u}\left(\frac{t}{t_{\mathrm{f}}}\right)^{n}\quad(u=x,y)\>,$ (40)
@@
$b_{8,u}$ $=$ $-315\quad,\quad b_{9,u}=70\quad(u=x,y)\>.$
The obtained solution for the STA dynamical-lattice trajectory – more precisely, the $x$ component $q_{0}^{\mathrm{x}}$ of the path of the potential minimum – is shown as a function of the transport time $t_{\mathrm{f}}$ in Fig. 4.
IV.2 eSTA dynamical-lattice trajectories
Having obtained STA dynamical-lattice trajectories in the problem at hand (cf. Sec. IV.1), we now proceed to find their STA counterparts based on the methodology reviewed in Sec. III.3. The control vectors corresponding to the sought-after eSTA trajectory will be denoted in what follows by $\bm{\zeta}_{\mathrm{x,S}}$ and $\bm{\zeta}_{\mathrm{y,S}}$; accordingly, this trajectory will be described by $q_{0}^{\mathrm{x}}(\bm{\zeta}_{\mathrm{x,S}};\>t)$ and $q_{0}^{\mathrm{y}}(\bm{\zeta}_{\mathrm{y,S}};\>t)$.
The eSTA paths of the potential minimum can be expressed as
$q_{0}^{\mathrm{x}}(\bm{\zeta}_{\mathrm{x,S}};\,t)$ $=$ $q_{0}^{\mathrm{x}}(\bm{\zeta}_{\mathrm{x},0};\,t)+f_{\mathrm{x}}(\bm{\alpha}_{\mathrm{x}};\,t)\>,$ (42)
$q_{0}^{\mathrm{y}}(\bm{\zeta}_{\mathrm{y,S}};\,t)$ $=$ $q_{0}^{\mathrm{y}}(\bm{\zeta}_{\mathrm{y},0};\,t)+f_{\mathrm{y}}(\bm{\alpha}_{\mathrm{y}};\,t)\>,$ (43)
where $\bm{\alpha}:=\left(\bm{\alpha}_{\mathrm{x}},\bm{\alpha}_{\mathrm{y}}\right)^{\intercal}$ is an auxiliary control vector. These eSTA paths are assumed to obey the conditions $q_{0}^{\mathrm{x}}(\bm{\zeta}_{\mathrm{x,S}};\,jt_{f}/7)=\zeta^{(j)}_{\mathrm{x},0}+\alpha^{(j)}_{\mathrm{x}}$ and $q_{0}^{\mathrm{y}}(\bm{\zeta}_{\mathrm{x,S}};\,jt_{f}/7)=\zeta^{(j)}_{\mathrm{y},0}+\alpha^{(j)}_{\mathrm{y}}$ ($j=1,\ldots,6$). The auxiliary functions $f_{u}(\bm{\alpha}_{u};\>t)$ ($u=x,y$) ought to satisfy the following boundary conditions:
$\begin{split}&f_{u}(\bm{\alpha}_{u};0)=f_{u}(\bm{\alpha}_{u};t_{f})=0\>,\\ &f_{u}(\bm{\alpha}_{u};jt_{f}/7)=\alpha^{(j)}_{u}\quad(\>j=1,\ldots,6\>)\>,\\ &\frac{\mathrm{d^{(n)}}}{\mathrm{d}t^{(n)}}f_{u}(\bm{\alpha}_{u};t^{\prime})|_{t^{\prime}=\{0,t_{f}\}}=0\quad(\>n=1,\ldots,4\>)\>.\end{split}$ (44)
These conditions are chosen such that $f_{u}(\bm{\alpha}_{u};\>t)$ can be controlled through $\bm{\alpha}_{u}$, satisfying at the same time the condition of continuity. Therefore, we choose the following Ansatz based on an eleventh-degree polynomial:
$f_{u}(\bm{\alpha}_{u};\>t)=\sum_{n=0}^{11}\sum_{j=1}^{6}\tilde{a}^{(j)}_{u,n}\alpha^{(j)}_{u}\left(\frac{t}{t_{f}}\right)^{n}\>.$ (45)
The concrete values of the coefficients $\tilde{a}^{(j)}_{\mathrm{x},n}$ and $\tilde{a}^{(j)}_{\mathrm{y},n}$ in these equations can be found in Table I of Hau .
[…]
@@ V Calculation of atomic dynamics
In what follows, we briefly present our chosen approach for evaluating the single-atom dynamics corresponding to STA- and eSTA-based dynamical-lattice trajectories [cf. Sec. IV]. This approach is based on the use of the FSOM for numerically solving the relevant TDSE (for a short introduction into this method, see Appendix C).
The crucial prerequisite for computing single-atom dynamics in the problem at hand amounts to evaluating the ground-state wave-function of the relevant optical-lattice potential. Within the general framework of the atom-transport problem (cf. Sec. II.1), that wave function represents the initial ($t=0$) atomic wave packet, i.e. the initial condition for single-atom transport. This wave function can be obtained to requisite accuracy using the plane-wave-expansion method (for details, see, e.g., Sto (b)), which entails Fourier expansions of the underlying lattice potential and the periodic part of single-particle Bloch wave functions in terms of reciprocal-lattice vectors. Yet, the same can be accomplished by employing the FSOM in conjunction with the ITE approach (for details, see Appendix D). Given that in this study we employ the FSOM for the dynamics part, this last scheme appears as the more convenient choice.
@@
where $\delta\dot{\mathbf{q}}_{0}(t)\equiv\dot{\mathbf{q}}_{0}(t+\delta t)-\dot{\mathbf{q}}_{0}(t)$. It is worthwhile to mention that two additional multiplicative terms – that do not depend on the momentum- and coordinate operators, hence leading to irrelevant time-dependent global phase factors – have been dropped on the right-hand-side (RHS) of the last equation, analogously to the dropped terms in Eq. (47) above.
With the aid of an analog of Eq. (C) and by performing a Fourier transformation of the $\hat{\mathbf{p}}-$dependent terms in (51), switching also to the coordinate representation of the momentum operator [i.e. $\hat{\mathbf{p}}\rightarrow(\hbar/i)\nabla$], we arrive at the FSOM-based (second-order) time-stepping scheme for the atom-transport problem at hand:
@@
In order to alleviate the computational burden in the problem at hand, we exploit the discrete translational symmetry of the system and restrict ourselves to the displacement of one unit cell of the original optical lattice. The appropriate choice of this unit cell has to be consistent with the transport direction. For instance, in order to evaluate transport in the $x$- and $y$ directions we make use of a rectangular unit cell (indicated by white dashed lines in Fig. 2). At the same time, describing transport in the diagonal direction necessitates the use of a $45$-degree rotated square-shaped unit cell (black dashed lines in Fig. 2).
@@
To carry out the spatial Fourier transformation using the FFT algorithm (cf. Appendix CNRc , we make use of a discrete two-dimensional $N_{\mathrm{x}}\times N_{\mathrm{y}}$ grid with $N_{\mathrm{x}}=N_{\mathrm{y}}=300$ points in the computational window whose dimensions (in units of $l_{\mathrm{x}}$ and $l_{\mathrm{y}}$) are $80$ and $80$, respectively. As regards the time-domain propagation of the TDSE in the comoving frame [cf. Eq. (47)], we carry it out using the adaptive approach NRc with a maximal relative error of $10^{-4}$, which could – strictly speaking – require additional (smaller) substeps in time. Yet, in all our runs, it proved sufficient to take $N_{t}$ between $20$ and $100$.